# Distance measurements from the internal dynamics of globular clusters: Application to the Sombrero galaxy (M 104)

### v1

@@ Abstract
Globular clusters (GCs) are overall ancient star clusters found in all massive galaxies (Brodie & Strader, 2006; Beasley, 2020). With stellar masses and compact sizes ( pc), they are dense enough to have survived for billions of years. Consequently, as some of the oldest stellar structures in the Universe, GCs have long been employed as ancient fossils of galaxy formation and assembly.
@@
In the Milky Way, spatially resolved studies have placed constraints on the chemical and kinematic properties of individual stars within GCs. For decades it has been known that GCs follow a tight relation between their magnitudes (or masses) and internal velocity dispersions (Pryor & Meylan, 1993). Similar to the Faber-Jackson relation for elliptical galaxies (Faber & Jackson, 1976), this relation is a projection of the fundamental plane that arises from the Virial Theorem (Djorgovski, 1995; Strader et al., 2009). However, as old stellar systems with a simple dynamical nature reflected in a narrow range of mass-to-light ratios (1 - 3; McLaughlin 2000; Baumgardt & Hilker 2018), the GC relation shows remarkably small scatter. As already discussed by Paturel & Garnier (1992), Djorgovski (1995), and Strader et al. (2009), this tight relation between GC magnitudes and velocity dispersion can be employed as an accurate distance indicator.
@@
Recently, Beasley et al. (2024) revisited the GC velocity dispersion (GCVD) distance relation by collecting -band magnitudes and velocity dispersions for GCs in the Milky Way and M 31 (Baumgardt & Hilker, 2018; Strader et al., 2011) in order to recalibrate the relation using geometric distances derived from Gaia parallaxes. By then applying the GCVD relation to Centaurus A using 57 GCs with velocity dispersion measurements from Dumont et al. (2022), Beasley et al. (2024) demonstrated that this method can yield distances with uncertainties as low as . Even more recently, the GCVD method was applied to derive a distance to the ultra diffuse galaxy NGC 1052-DF2, achieving an uncertainty of based on 5 GCs (Beasley et al., 2025).
@@
As of now, application of the GCVD distance method is limited by the availability of velocity dispersion measurements. Outside the Local Group, dispersion measurements are rare as they rely on moderately high spectral resolution () to measure dispersions down to km s-1, a typical value for a GC with a mass of (the approximate mean mass of GCs in the Milky Way, Baumgardt et al. 2020). In an ongoing effort to collect larger samples of GC velocity dispersions, we present here velocity dispersions of 85 GCs and one ultra-compact dwarf galaxy (UCD) around M 104 (NGC 4594, the Sombrero galaxy), based on high-resolution spectroscopy with the FLAMES GIRAFFE multi-object spectrograph at the Very Large Telescope (VLT). In Fahrion et al. (2025), we presented an extensive catalogue of GC velocities using the same dataset combined with spectra from the OSIRIS instrument at the Gran Telescopio Canarias (GTC) and analysis of archival VLT MUSE data. In this paper, our focus is on deriving a GCVD distance to M 104. Additionally, for a subsample of 29 GCs with available size measurements, we derive dynamical masses and mass-to-light ratios.
@@
M 104 is a particularly interesting target for the GCVD method. As a peculiar galaxy with its prominent, almost edge-on disc, M 104 has been studied extensively using space- and ground-based facilities (e.g. Emsellem 1995; Harris et al. 2010; Jardel et al. 2011; Gadotti & Sánchez-Janssen 2012; Cohen et al. 2020). Also, its rich GC system has been the focus of many studies (e.g. Wakamatsu 1977; Harris et al. 1984; Bridges & Hanes 1992; Bridges et al. 1997; Forbes et al. 1997; Larsen et al. 2002; Rhode & Zepf 2004; Bridges et al. 2007; Alves-Brito et al. 2011; Dowell et al. 2014; Kang et al. 2022). The total number of GCs is estimated to lie between 1500 and 2000 (Rhode & Zepf, 2004; Kang et al., 2022), and combining literature samples with newly collected data, we collected line-of-sight velocities for 499 of those in Fahrion et al. (2025), corresponding to approximately 25% of the total GC system. With such a rich GC system, M 104 is therefore an ideal target to apply the GCVD method and test it against other distance measurements.
@@
Distance measurements to M 104 from the literature range between 7 and 22 Mpc111From the NASA/IPAC Extragalactic Database and encompass a variety of different distance indicators. Using the tip of the red giant branch (TRGB) in deep Hubble Space Telescope (HST) observations, McQuinn et al. (2016) derived a distance of (sys.) Mpc, which is the most recent precise distance measurement to M 104. Older distance measurements using surface brightness fluctuations (SBF) based on infrared HST observations, Jensen et al. (2003) reported a distance of 9.08 Mpc, similar to other SBF measurements (Ciardullo et al., 1993; Ferrarese et al., 2000; Tonry et al., 2001) that found distances between 8 and 10 Mpc. Similar values were also reported based on the luminosity functions of GCs (GCLF; Bridges & Hanes 1992; Spitler et al. 2006) and planetary nebulae (PNLF; Ciardullo et al. 1993; Ford et al. 1996).
@@
The aim of this paper is to construct the GCVD relation for M 104 GCs for a precise distance measurement. The paper is structured as follows: In Sect. 2, we describe the FLAMES high-resolution data. Sect. 3 describes our methods of measuring dispersions and correcting for aperture effects. The distance measurement is detailed in Sect. 5 and Sect. 4 describes the mass-to-light ratios for a subsample of GCs. We summarise and conclude in Sect. 7.
@@ 2 Data
The data were reduced using standard procedures using the FLAMES GIRAFFE esorex pipeline and custom scripts, with the variation that the sky-subtraction was performed using pPXF (Cappellari & Emsellem, 2004; Cappellari, 2017, 2023) to fit each spectrum with a stellar model from the PHOENIX library (Husser et al., 2013) and sky templates as sky input. After sky subtraction, spectra from individual exposures were corrected for barycentric velocities, resampled, and combined using a median weighted by the signal-to-noise (S/N) ratio of each spectrum.
@@ 3.1 Spectral fitting
We fit the FLAMES GC spectra with pPXF (Cappellari & Emsellem, 2004; Cappellari, 2017, 2023). pPXF fits the spectrum with a library of suitable model or empirical spectra and returns the parameters of the line-of-sight velocity distribution (LOSVD) such as radial velocity and velocity dispersion. We chose additive polynomials of degree 4 and multiplicative polynomials of degree 1 to account for variations in the continuum and flux calibrations, but as described in Beasley et al. (2025), the results are not very sensitive to the choice of degree of the polynomials. Because sky residuals remained in the reduced data, we carefully masked those lines using a custom mask for every spectrum. We began by normalising the spectrum without sky subtraction in a continuum wavelength range to the continuum level of the sky-subtracted spectrum. Then, we masked out all wavelength pixels in which the normalised sky spectrum exceeds 1.2 times the continuum level. We fitted all spectra in a wavelength range between 8450 and 8850 Å except for four GCs that had strong sky residuals at redder wavelengths. Here, we restricted the upper wavelength range to 8750 Å due to strong sky residuals at longer wavelengths.
We used the high-resolution stellar population models from Coelho (2014) to fit the FLAMES spectra. In a first step, we selected red (sub) giant stars with effective temperature between 4500 and 5000 K, surface gravity between log(g)= 2.0 and 4.0, and metallicities between 1.5 and 0 dex. Then, the selected model spectra were brought to a spectral resolution of 0.54 Å to match the FLAMES spectra. Figure 1 shows eight GC spectra and their fits as examples.
@@ 3.2 Aperture Corrections
We give the full sample of velocity dispersion measurements in Appendix A. The values range between 4 and 30 km s-1 with mean uncertainties of 2.5 km s-1. For SUCD1, we measured an aperture-corrected dispersion of km s-1, in agreement with the measurement by Hau et al. (2009) of km s-1 based on a lower resolution Keck/DEIMOS spectrum.
@@ 3.3 Available photometry
Moreover, 68 GCs have -band magnitudes available from Rhode & Zepf (2004) based on imaging with the Mosaic detector on the Kitt Peak 4 metre telescope. The ones with -band magnitudes available from both studies show good agreement within 0.1 mag. In total, -band magnitudes are available for 77 of 85 suitable GCs, in a magnitude range of mag. For SUCD1, Hau et al. (2009) reported a -band magnitude of 17.46 mag based on the same HST/ACS observations as used by Spitler et al. (2006) and Harris et al. (2010). In addition, all GCs have -band magnitudes from Kang et al. (2022) available, based on CFHT/MegaCam imaging.
@@
4 Mass-to-light Ratios
For a subsample of 29 GCs and SUCD1, size measurements are available based on HST imaging (Hau et al., 2009; Harris et al., 2010). We derive -band mass-to-light ratios () using the aperture-corrected global velocity dispersions () and half-light radii (). We obtain virial masses for the Sombrero GCs from Eq. 1 (e.g. Larsen et al. 2002; Beasley et al. 2025):
@@ 4 Mass-to-light Ratios
The derived mass-to-light ratios are shown in Fig. 4 compared to GCs around Centaurus A (Dumont et al., 2022), M 31 (Strader et al., 2011), and the Milky Way (Baumgardt et al., 2020). The values are colour-coded according to their metallicities, derived from the FLAMES spectra (see Fahrion et al. 2025). We also included a line marking the mean of the MW GCs (, Beasley et al. 2025 based on data from Baumgardt et al. 2020).
We find values in a range from , with a mean for this sample of M 104 GCs. Many of them have higher mass-to-light ratios than the MW population, but they are also considerably brighter than most of the Galactic GCs. The brightest M 104 GCs are even significantly brighter than Centauri, the most massive GC of the Milky Way, that also shows an increased mass-to-light ratio.
We find reasonable agreement with GCs in Centaurus A (), but a direct comparison is difficult based on the low number of size measurements available for our M 104 sample. Similar to the behaviour seen in the Centaurus A GCs (Dumont et al., 2022), we also observe a range in mass-to-light ratio that exceeds those of the Milky Way GCs (M/L 1.5). From their analysis, Dumont et al. (2022) interpreted this scatter with a population of stripped nuclei masquerading as GCs. In M 104, we find that SUCD1 indeed shows an elevated mass-to-light ratio, in line with other UCDs (e.g. Rejkuba et al. 2007; Mieske et al. 2008). Nonetheless, the defining feature that sets it apart from the GC population is its significantly larger luminosity, being about 1.4 magnitudes brighter than the next brightest GC. Based on its magnitude, size, and velocity dispersion, we derive a dynamical mass of , similar to the value reported by Hau et al. (2009).
In the lower panel of Fig. 4, we also included GCs that lack direct size measurements. For those, we assumed a size of 2.44 pc, corresponding to the median of size measurements as reported in Harris et al. (2010). These estimated values show a similar range, with several GCs having similar values as Milky Way GCs. However, we note that the uncertainties increase substantially for the fainter GCs.

### v2

# Distance measurements from the internal dynamics of globular clusters: Application to the Sombrero galaxy (M104)

@@ Abstract
Globular clusters (GCs) are ancient star clusters found in all massive galaxies (Brodie and Strader, 2006; Beasley, 2020). With stellar masses and compact sizes ( pc), they are dense enough to have survived for billions of years. Consequently, as some of the oldest stellar structures in the Universe, GCs have long been employed as ancient fossils of galaxy formation and assembly.
@@
In the Milky Way, spatially resolved studies have placed constraints on the chemical and kinematic properties of individual stars within GCs. For decades it has been known that GCs follow a tight relation between their magnitudes (or masses) and internal velocity dispersions (Pryor and Meylan, 1993). Similar to the Faber-Jackson relation for elliptical galaxies (Faber and Jackson, 1976), this relation is a projection of the fundamental plane that arises from the Virial Theorem (Djorgovski, 1995; Strader et al., 2009): In a self-gravitating system, the total mass scales with , where refers to the half-light radius. As old stellar systems with a narrow size distribution (pc Masters et al. 2010) and a narrow range of mass-to-light ratios (1 - 3; McLaughlin 2000; Baumgardt and Hilker 2018), the GC relation shows remarkably small scatter. As already discussed by Paturel and Garnier (1992), Djorgovski (1995), and Strader et al. (2009), this tight relation between GC magnitudes and velocity dispersion can be employed as an accurate distance indicator.
@@
Recently, Beasley et al. (2024) revisited the GC velocity dispersion (GCVD) distance relation by collecting -band magnitudes and velocity dispersions for GCs in the Milky Way and M31 (Baumgardt and Hilker, 2018; Strader et al., 2011) in order to recalibrate the relation using geometric distances derived from Gaia parallaxes. By then applying the GCVD relation to Centaurus A using 57 GCs with velocity dispersion measurements from Dumont et al. (2022), Beasley et al. (2024) demonstrated that this method can yield distances with uncertainties as low as . Even more recently, the GCVD method was applied to derive a distance to the ultra diffuse galaxy NGC 1052-DF2, achieving an uncertainty of based on 5 GCs (Beasley et al., 2025).
@@
As of now, application of the GCVD distance method is limited by the availability of velocity dispersion measurements. Outside the Local Group, dispersion measurements are rare as they rely on moderately high spectral resolution () to measure dispersions down to km s-1, a typical value for a GC with a mass of (the approximate mean mass of GCs in the Milky Way, Baumgardt et al. 2020). In an ongoing effort to collect larger samples of GC velocity dispersions, we present here velocity dispersions of 85 GCs and one ultra-compact dwarf galaxy (UCD) around M104 (NGC 4594, the Sombrero galaxy), based on high-resolution spectroscopy with the FLAMES GIRAFFE multi-object spectrograph at the Very Large Telescope (VLT). In Fahrion et al. (2025), we presented an extensive catalogue of GC velocities using the same dataset combined with spectra from the OSIRIS instrument at the Gran Telescopio Canarias (GTC) and analysis of archival VLT MUSE data. In this paper, our focus is on deriving a GCVD distance to M104. Additionally, for a subsample of 29 GCs with available size measurements, we derive dynamical masses and mass-to-light ratios, allowing us to test the effect of sizes and mass-to-light ratios on the plane.
@@
M104 is a particularly interesting target for the GCVD method. As a peculiar galaxy with its prominent, almost edge-on disc, M104 has been studied extensively using space- and ground-based facilities (e.g. Emsellem 1995; Harris et al. 2010; Jardel et al. 2011; Gadotti and Sánchez-Janssen 2012; Cohen et al. 2020). Also, its rich GC system has been the focus of many studies (e.g. Wakamatsu 1977; Harris et al. 1984; Bridges and Hanes 1992; Bridges et al. 1997; Forbes et al. 1997; Larsen et al. 2002; Rhode and Zepf 2004; Bridges et al. 2007; Alves-Brito et al. 2011; Dowell et al. 2014; Kang et al. 2022). The total number of GCs is estimated to lie between 1500 and 2000 (Rhode and Zepf, 2004; Kang et al., 2022), and combining literature samples with newly collected data, we collected line-of-sight velocities for 499 of those in Fahrion et al. (2025), corresponding to approximately 25% of the total GC system. With such a rich GC system, M104 is therefore an ideal target to apply the GCVD method and test it against other distance measurements.
@@
Distance measurements to M104 from the literature range between 7 and 11 Mpc111From the NASA/IPAC Extragalactic Database. Additional estimates using the Tully-Fisher relation give generally larger distances ( 11 - 22 Mpc; Bottinelli et al. 1984; Tully and Fisher 1988; Sorce et al. 2014), but it is questionable if this relation can be applied to M104 with its peculiar morphology and gas distribution (see e.g. McQuinn et al. 2016). and encompass a variety of different distance indicators. Using the tip of the red giant branch (TRGB) in deep Hubble Space Telescope (HST) observations, McQuinn et al. (2016) derived a distance of (sys.) Mpc, which is the most recent precise distance measurement to M104. Older distance measurements using surface brightness fluctuations (SBF) based on infrared HST observations, Jensen et al. (2003) reported a distance of 9.08 Mpc, similar to other SBF measurements (Ciardullo et al., 1993; Ferrarese et al., 2000; Tonry et al., 2001) that found distances between 8 and 10 Mpc. Similar values were also reported based on the luminosity functions of GCs (GCLF; Bridges and Hanes 1992; Spitler et al. 2006) and planetary nebulae (PNLF; Ciardullo et al. 1993; Ford et al. 1996).
@@
The aim of this paper is to construct the GCVD relation for M104 GCs for a precise distance measurement. The paper is structured as follows: In Sect. 2, we describe the FLAMES high-resolution data. Sect. 3 describes our methods of measuring dispersions, the available photometry and our aperture correction. Sect. 4 describes the distribution of M104 GCs on the - plane. The distance measurement is detailed in Sect. 5 and compared to literature distances in Sect. 6. We summarise and conclude in Sect. 7.
@@ 2 Data
The data were reduced using standard procedures using the FLAMES GIRAFFE esorex pipeline and custom scripts, with the variation that the sky-subtraction was performed using pPXF (Cappellari and Emsellem, 2004; Cappellari, 2017, 2023) to fit each spectrum with a stellar model from the PHOENIX library (Husser et al., 2013) and sky templates as sky input. After sky subtraction, spectra from individual exposures were corrected for barycentric velocities, resampled, and combined using a median weighted by the signal-to-noise (S/N) ratio of each spectrum.
@@ 3.1 Spectral fitting
We fit the FLAMES GC spectra with pPXF (Cappellari and Emsellem, 2004; Cappellari, 2017, 2023). pPXF fits the spectrum with a library of suitable model or empirical spectra and returns the parameters of the line-of-sight velocity distribution (LOSVD) such as radial velocity and velocity dispersion. We chose additive polynomials of degree 4 and multiplicative polynomials of degree 1 to account for variations in the continuum and flux calibrations, but as described in Beasley et al. (2025), the results are not very sensitive to the choice of degree of the polynomials. Because sky residuals remained in the reduced data, we carefully masked those lines using a custom mask for every spectrum. We began by normalising the spectrum without sky subtraction in a continuum wavelength range to the continuum level of the sky-subtracted spectrum. Then, we masked out all wavelength pixels in which the normalised sky spectrum exceeds 1.2 times the continuum level. We fitted all spectra in a wavelength range between 8450 and 8850 Å except for four GCs that had strong sky residuals at redder wavelengths. Here, we restricted the upper wavelength range to 8750 Å due to strong sky residuals at longer wavelengths.
We used the high-resolution stellar population models from Coelho (2014) to fit the FLAMES spectra. In a first step, we selected red (sub) giant stars with effective temperature between 4500 and 5000 K, surface gravity between log(g)= 2.0 and 4.0, and metallicities between 1.5 and 0 dex. Then, the selected model spectra were brought to a spectral resolution of 0.54 Å to match the FLAMES spectra. Figure 1 shows eight GC spectra and their fits as examples, chosen to cover a range of -band magnitudes and signal-to-noise ratios.
@@ 3.2 Aperture Corrections
[…]
We give the full sample of velocity dispersion measurements in Appendix B. The values range between 4 and 30 km s-1 with mean uncertainties of 2.5 km s-1. For SUCD1, we measured an aperture-corrected dispersion of km s-1, in agreement with the measurement by Hau et al. (2009) of km s-1 based on a lower resolution Keck/DEIMOS spectrum.
@@ 3.3 Available photometry
Moreover, 68 GCs have -band magnitudes available from Rhode and Zepf (2004) based on imaging with the Mosaic detector on the Kitt Peak 4 metre telescope. The ones with -band magnitudes available from both studies show good agreement within 0.1 mag. In total, -band magnitudes are available for 77 of 85 suitable GCs, in a magnitude range of mag. For SUCD1, Hau et al. (2009) reported a -band magnitude of 17.46 mag based on the same HST/ACS observations as used by Spitler et al. (2006) and Harris et al. (2010). In addition, all GCs have -band magnitudes from Kang et al. (2022) available, based on CFHT/MegaCam imaging.
@@
4 The Sombrero GCs on the - plane
Before deriving a GCVD distance, we explore the properties of the GC sample in the - plane to test underlying dependencies with GC parameters. In Fig. 4, we colour-coded the GCs according to different parameters to test if the scatter correlates with GC parameters. For the full sample, we collected -band magnitudes from the HST study by Harris et al. (2010) and ground-based measurements from Rhode and Zepf (2004).
4.1 Comparison to other GC samples
The top left panel of Fig. 4 compares the band magnitudes and velocity dispersions of the M104 GCs with that of the Milky Way and M31 (white circles), which were used to calibrate the GCVD relation (Beasley et al., 2024), assuming a distance of 9 Mpc. Additionally, GCs in Centaurus A are shown with orange triangles (Dumont et al., 2022). As this panel shows, the M104 GCs cover a very similar range in magnitudes and dispersions as the Cen A GCs. The scatter in each direction appears similar to the calibration sample. SUCD1 is about a magnitude brighter than even the brightest GC in the calibration sample or Cen A.
4.2 Metallicities, colours and velocities
In the second panel on the top left of Fig. 4, the GCs are colour-coded according to their metallicity. The metallicity values are drawn from the full M104 sample described in Fahrion et al. (2025). Where available, we used metallicities from optical VLT/MUSE or GTC/OSIRIS spectroscopy, and for the remaining GCs the metallicities were derived directly from full spectrum fitting to the FLAMES spectra. While the GCs with measured dispersions cover a broad range of metallicities, there seems to be no clear correlation with the scatter in the magnitude-dispersion plane. Splitting the MW and M31 GCs into metal-poor and metal-rich GCs, Beasley et al. (2024) found no difference in the GCVD relation’s slope or offset, in agreement with our finding for M104.
4.3 Projected distances
The bottom left panel in Fig. 4 shows the GCs colour-coded by their projected distance to the centre of M104. The projected distances range from 1 to 20, and no obvious relation with the scatter in the - plane is apparent. We note that the brightest GCs in our sample are also relatively close (a few arcminutes) to the centre. This could be a selection effect or could reflect mass segregation of the GC sample since more massive GCs tend to experience stronger dynamical friction effects (e.g. Tremaine 1976; Turner et al. 2012; Lotz et al. 2001). In the Appendix A we show the magnitudes and dispersions as a function of galactocentric distance.
4.4 Globular cluster sizes
As the GCVD relation is a consequence of the Virial Theorem, based on the relation between total mass and the combination of velocity dispersion squared and size (), the size influences the position of a source on the relation, with the effect expected to scale linearly. This is seen in the second panel in the bottom row of Fig. 4, where smaller-sized GCs lie to the right of the GCVD relation. Also, the observed offset of SUCD1 could be explained by the size difference as it is a particularly large object with a reported half-light radius of pc (Hau et al., 2009), significantly larger than typical GCs (mean pc, Harris et al. 2010; Masters et al. 2010).
Therefore, to test if the GCVD relation is applicable to M104, we compare the sizes of its GCs with the Milky Way (from Baumgardt and Hilker 2018) and M31 GCs (from Barmby et al. 2007; Peacock et al. 2010; Strader et al. 2011) that were used for the original calibration of the relation (Beasley et al., 2024). Figure 5 shows this comparison, based on the size measurements of 652 M104 GCs from Harris et al. (2010) which were based on HST ACS imaging. We find that the size distribution of M104 GCs aligns well with the sizes of the calibration MW and M31 sample.
However, we note that the small subsample of 29 GCs for which we have both sizes and dispersion measurements (shown in pink in Fig. 5) contains preferentially larger GCs. As these GCs are also among the brighter GCs in our sample (see Fig. 4), we believe that this is reflecting a slight brightness-size relation in this small subsample (Harris et al., 2010). The larger sample of GCs with dispersion measurements likely shows a similar size distribution as the calibration sample, but more size measurements would be needed to confirm this.
4.5 Mass-to-light ratios
The GCVD distance indicator relies on the assumption that all GCs follow a similar relation between observed light and their mass. To test this assumption, we derived dynamical mass-to-light ratios as described the -band mass-to-light ratios () using the aperture-corrected global velocity dispersions () and half-light radii () from Harris et al. (2010). We obtain virial masses for the Sombrero GCs from Eq. 1 (e.g. Larsen et al. 2002; Beasley et al. 2025):
@@ 4.5 Mass-to-light ratios
The derived mass-to-light ratios are shown in Fig. 6 compared to GCs around Centaurus A (Dumont et al., 2022), M31 (Strader et al., 2011), and the Milky Way (Baumgardt et al., 2020). We show here both the GCs for which size measurements are available (larger circles) and we also included GCs that lack direct size measurements (smaller circles). For those, we assumed a size of 2.44 pc, corresponding to the median of size measurements as reported in Harris et al. (2010). These estimated values show a similar range, with several GCs having similar values as Milky Way GCs. However, we note that the uncertainties increase substantially for the fainter GCs. For M104 GCs, we find values in a range from , with a mean and a dispersion around the mean of . Many M104 GCs have higher mass-to-light ratios than the MW population, but they are also considerably brighter than most of the Galactic GCs. The brightest M104 GCs are even significantly brighter than Centauri, the most massive GC of the Milky Way, that also shows an increased mass-to-light ratio.
We find reasonable agreement with GCs in Centaurus A (), but a direct comparison is difficult based on the low number of size measurements available for our M104 sample. Similar to the behaviour seen in the Centaurus A GCs (Dumont et al., 2022), we also observe a range in mass-to-light ratio that exceeds those of the Milky Way GCs. From their analysis, Dumont et al. (2022) interpreted this scatter with a population of stripped nuclei masquerading as GCs. In M104, we find that SUCD1 indeed shows an elevated mass-to-light ratio, in line with other UCDs (e.g. Rejkuba et al. 2007; Mieske et al. 2008). Nonetheless, the defining feature that sets it apart from the GC population is its significantly larger luminosity, being about 1.4 magnitudes brighter than the next brightest GC. Based on its magnitude, size, and velocity dispersion, we derive a dynamical mass of , similar to the value reported by Hau et al. (2009).
The second panel in the bottom row of Fig. 4 shows the - plane colour-coded by the -band mass-to-light ratios. GCs with larger mass-to-light ratios appear to lie preferentially to the right of the GCVD relation. Given that the dynamical mass scales with the square of the velocity dispersion, this is perhaps not surprising, but indicates that the GCVD distance method might only by applicable for objects within a narrow range of mass-to-light ratios. Interestingly, SUCD1 is an outlier in this respect as it lies above the relation, but still has a larger mass-to-light ratio than most of the GCs. This is likely a consequence of the elevated mass-to-light ratio of SUCD1 stemming from its large size, while the higher mass-to-light ratios of the GCs are driven by an increased velocity dispersion.
4.6 Relaxation times
The final panel in Fig. 4 shows the scatter in relation to the half-mass relaxation times , estimated from the sizes and dynamical masses of the GCs. To obtain a measure of , we followed Binney and Tremaine (1987):
where ln() is used to describe the Coulomb logarithm, with as the number of stars within the GC, the average mass of a star, the mass of the GC, and the half-light radius. We assume (Giersz and Heggie, 1994), and . We find relaxation times ranging from 0.6 to 10 Gyr, while the relaxation time for SUCD1 is estimated at Gyr due to its large mass.
As the relaxation time depends on both the mass and size of the GCs, it is not surprising that the scatter in the GCVD relation shows a similar trend as with the half-light radii: smaller GCs have lower relaxation times and lie to the right of the relation.